OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 15 and a15 = –41 ?

OpenStudy (zarkon):

-145

OpenStudy (anonymous):

15-1=14 iterations of the sequence. Delta(a)=-41-15=-56 15-(41-1)4=15-160=-145

OpenStudy (anonymous):

Using the same procedure I gave you in the last question you can answer this one